an executive leaves the plant and travels on a highway at 55mi/h to headquarters. he returns by helicopter on a route parallel to the highway at 125 mi/h. the total travel time is 1.8 h. how far is it from the plant to headquarters?
Answer by ankor@dixie-net.com(12698)

Travel_Word_Problems/52735 (2006-09-18 16:10:01)
:
an executive leaves the plant and travels on a highway at 55 mi/h to headquarters. he returns by helicopter on a route parallel to the highway at 125 mi/h. the total travel time is 1.8 h. how far is it from the plant to headquarters?
Assuming the drive dist & the helicopter dist are the same:
:
Find the time for driving, then it's simple to find the distance.
:
Let t = drive time
Then [1.8-t] = flying time
:
Dist = speed * time
:
drive dist = fly dist
55 * t = 125 * [1.8-t]
55t = 225 - 125t
55t + 125t = 225
180t = 225
t = 225/180
t = 1.25 hrs to drive
:
Find dist using drive time: 1.25 * 55 = 68.75
:
Fly time = 1.8 - 1.25 = .55 hrs
Check dist with fly time: .55 * 125 = 68.75 mi also